How to calculate equilibrium income using the multiplier???



Mpc=0.75



Multiplier =4



Investment =200 million



Autonomous consumption =100 million

1
Expert's answer
2022-01-14T14:44:08-0500

Solution:

Equilibrium income = Multiplier x AD

AD = C + I

C = 100 + 0.75Y

Y = 100 + 0.75Y + 200

Y – 0.75Y = 300

0.25Y = 300

Y = 1,200

Equilibrium income = 1,200 "\times" 4 = 4,800

Equilibrium income = 4,800